# Fenchel-Young Variational Inference

### v2

@@ Abstract
Bayesian learning and inference seek distributions of the unknowns—parameters and/or latent variables—that balance empirical risk and regularization. In both standard and variational Bayesian inference, this disagreement is measured by the Kullback-Leibler divergence (KLD). We present Fenchel-Young variational inference (FYVI), which broadens previous formulations by replacing the KLD with a Fenchel-Young (FY) regularizer. We focus on a class of FY losses based on Tsallis entropies that contains the standard KLD as a particular case, and which allows us to obtain generalized variational posterior distributions with support strictly smaller than that of the prior. FYVI includes new generalized notions of free energy, evidence, and posterior. We derive alternating minimization and gradient backpropagation algorithms to approximate the generalized FY posterior. Our FY expectation-maximization algorithm and FY latent variable models can be seen as generalized variants of classical algorithms and models. The proposed methods are empirically competitive, outperforming established counterparts in controlled experiments, in which directly modeling limited support (sparse) posteriors is intuitively appealing.
@@ 1 Introduction
Much of machine learning can be formulated as variational optimization problems aiming to minimize a regularized empirical risk (average loss) over a training set, where the regularizer encourages the solution—learned parameters—to have certain properties. In Bayesian approaches to machine learning and inference, all the unknowns—learnable parameters and, possibly, latent/hidden variables—are seen as random variables, and learning becomes synonymous with inferring the conditional distribution of the unknowns given the observed data. Based on the variational formulation of Bayesian inference by Zellner [59] and the generalized loss posteriors of Bissiri et al. [7], Khan and Rue [32] unify diverse, well-established Bayesian inference algorithms under a common formal umbrella: finding the (generalized) posterior distribution (in some allowed family) of the unknown parameters or latent variables that strikes a balance between a low expected risk and a small Kullback-Leibler divergence (KLD) from a given prior.
In this paper, we extend the variational formulations of (generalized) posterior inference studied along related lines by Bissiri et al. [7], Knoblauch et al. [34], and Khan and Rue [32], by considering alternative optimality criteria that exploit the class of Fenchel-Young (FY) losses [10, 39]. In particular, we focus on FY losses based on Tsallis entropies, which subsume the Kullback-Leibler divergence as a particular case. We generalize the formulations of Knoblauch et al. [34] and Khan and Rue [32], leading to a strictly wider class of statistical families, which includes generalized posteriors with limited (non-full, a.k.a. sparse) support, as well as heavy-tailed posteriors. [^1] In this paper, we focus on the less studied case of sparse posteriors, rather than on heavy-tailed ones, which are a non-trivial case that also falls under the framework we propose.
@@
As reviewed in the next section, Zellner [59] rediscovers the fact that the Bayesian posterior is the solution of an unconstrained optimization/variational problem over the space of distributions where the objective function(al) is the expected log-likelihood plus the Kullback-Leibler divergence (KLD) from the prior [20, 16]. The KLD term in the variational formulation of posterior inference is equivalent to a maximum entropy regularizer that favors posteriors with high Shannon entropy [31], giving a privileged position to exponential families in this context [25], following from the well-known fact that these are maximum Shannon entropy distributions, under expectation constraints.
If instead of the KLD, we use FY losses induced by Tsallis entropies, the corresponding generalized posteriors belong to deformed exponential families, which, unlike standard exponential families, may have variable support. Whereas Knoblauch et al. [34] analyze arbitrary combinations of divergences, losses, and exponential variational families, we show that when using FY losses based on Tsallis entropies, the variational family cannot be chosen independently of the divergence; in fact, it is determined by it.
In §3 we propose the Fenchel-Young variational infrence (FYVI) framework, which includes new generalized notions of FY free energy, FY evidence, FY evidence lower bound, and FY posterior. We derive an EM-style [19] algorithm for parameter estimation, and describe gradient backpropagation for amortized inference with deformed exponential family posteriors. Both of these estimation methods can be used to fit arbitrary generative models that may be specified as instances of FYVI.
@@
In §4 we present experiments with the FYEM algorithm applied to Gaussian mixture model (GMM) [42] estimation, and two FY variational autoencoder models (VAEs) [33, 50], one applied to images and the other to documents. The FYEM algorithm for GMMs recovers the classical EM as a special case and yields a new generalization with an adaptively sparse E-step. In the image FYVAE we obtain the best performance when we combine a sparse posterior with a sparse observation model. We also outperform the VAE baseline of a neural variational document model (NVDM) [40], using a FY loss in place of a negative log-likelihood combined with a standard Gaussian latent variable, by tuning the sparsity of the observation model.
Our code will be publicly released upon acceptance.
@@ 2.2 Variational View of Bayesian Inference
2.2.1 Standard Bayesian Inference
As has been shown [12, 16, 20, 59], the Bayesian posterior $p_{Z|X}$ can be expressed in variational form as the distribution in $\Delta(\mathcal{Z})$ which strikes a balance between having a small KLD from the prior $p_{Z}$, and yielding a high expected log-likelihood $\log p_{X|Z}$, given some observation $x\in\mathcal{X}$. Formally,
${p_{Z|X}(\cdot|x)}$ $=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\underset{Z\sim q}{\mathbb{E}}\bigl{[}\log p_{X|Z}(x|Z)\bigr{]}\!-\!D_{\textnormal{KL}}(q\|p_{Z})$ (1)
$=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\underset{Z\sim q}{\mathbb{E}}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}\!+\!H(q).$ (2)
The optimum of (1), achieved by setting $q=p_{Z|X}(\cdot|x)$, is (after trivial manipulation) equal to the log-evidence:
$\mathbb{E}_{Z\sim p_{Z|X}(\cdot|x)}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}+H\bigl{(}p_{Z|X}(\cdot|x)\bigr{)}=\log p_{X}(x).$ (3)
More often than not, the posterior and evidence cannot be easily obtained, which has stimulated decades of work on approximate Bayesian methods.
2.2.2 Variational Inference
In variational inference (VI) [9, 31], the goal is to approximate the posterior and bound the evidence when these are intractable, by confining $q$ to a tractable family, i.e. a subset of the entire space of probability functions, $\mathcal{Q}\subseteq\Delta(\mathcal{Z})$. The familiar VI optimization problem is
@@ 2.2.2 Variational Inference
and its optimal value is the ELBO (evidence lower bound): ${\mathbb{E}_{Z\sim q_{x}^{\star}}}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}+H\bigl{(}q_{x}^{\star}\bigr{)}\leq\log p_{X}(x)$. The negative of the objective function in (4) is often called the variational free energy.
2.2.3 Generalized Variational Inference
@@
Khan and Rue [32] use Bissiri et al. [7]’s extension of the classical variational problem in (4) to generalized posteriors, in order to derive as special cases much of the core catalog of modern machine learning algorithms and to propose new ones. This extension consists of replacing the negative log-likelihood with an arbitrary loss function $\ell(x;Z)$:
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)\right]+D_{\text{\footnotesize KL}}(q\|p_{Z})=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)-\log p_{Z}(Z)\right]-H(q),$ (5)
@@
where the notation $q_{x}^{\star}$ highlights the dependency on the observations $x$. Khan and Rue [32] restrict $\mathcal{Q}$ to be a regular minimal exponential family, i.e., $\mathcal{Q}=\{q(z)=h(z)\exp\bigl{(}\langle\lambda,t(z)\rangle-A(\lambda)\bigr{)},\mbox{for }\lambda\in\Lambda\}$, where $\Lambda$ is the parameter space, $t$ is a canonical sufficient statistic, $h$ is the base measure (independent of $\lambda$), and $A(\lambda)$ is the log-partition function. In a similar vein, Knoblauch et al. [34] propose generalized VI (GVI), by additionally considering divergences other than KLD, and introducing the rule of three (RoT), which refers to the choice of the three elements required for GVI: a loss function, a divergence measuring deviation of variational posterior from prior, and a set of allowed solutions $\mathcal{Q}$.
2.3 Fenchel-Young Losses and Related Work
@@
In a previously unrelated line of research, Blondel et al. [10] propose Fenchel-Young (FY) losses as a generalization of the ubiquitous cross-entropy loss that can be seen as mixed space variants of Bregman divergences [4], [10, §3.2]. FY losses have been used for supervised classification problems, but not in any kind of (generalized) Bayesian contexts, such as VI or latent variable models. Our notation below treats both discrete and continuous domains in a unified way.
@@
Let $\Omega:\Delta(\mathcal{Z})\rightarrow\bar{\mathbb{R}}$ be a lower semi-continuous (l.s.c.) proper convex function, and let $\Omega^{*}$ be its Fenchel-Legendre conjugate [^2] Given a function $f:\mathcal{U}\rightarrow\bar{\mathbb{R}}$, its Fenchel conjugate $f^{*}:\mathcal{U}^{*}\rightarrow\bar{\mathbb{R}}$ is defined as $f^{*}(x)=\sup_{u\in\mathcal{U}}(\langle x,u\rangle-f(u)),$ where $\mathcal{U}^{*}$ is the dual space of $\mathcal{U}$. If $f$ is convex l.s.c., biconjugation holds: $f^{**}=f$. If $\mathcal{U}$ is compact or $f$ is continuous, sup may be replaced with max, because the supremum is achieved. In this case, assuming differentiability, the maximizing argument in the definition of $f^{*}$ is $\arg\max_{u\in\mathcal{U}}(\langle x,u\rangle-f(u))=\nabla f^{*}(x)$. [11, 6]. Consider also a score function $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ that assigns a real-valued score to each element of $\mathcal{Z}$. In a Bayesian setting, $\eta$ corresponds to a log-prior. For a distribution $q\in\Delta(\mathcal{Z})$, the FY loss induced by $\Omega$ is a measure of how incompatible $\eta$ and $q$ are:
$L_{\Omega}(\eta;q)=\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)=\Omega^{*}(\eta)-\mathbb{E}_{Z\sim q}[\eta(Z)]+\Omega(q),$ (6)
@@
where the second equality follows from the simple fact that $\mathbb{E}_{Z\sim q}[\eta(Z)]=\langle q,\eta\rangle$, for discrete or continuous $\mathcal{Z}$. An FY loss is always non-negative, due to the Fenchel-Young inequality [6],
$\Omega^{*}(\eta)=\max_{q\in\Delta(\mathcal{Z})}\bigl{(}\langle q,\eta\rangle-\Omega(q)\bigr{)}\;\Rightarrow\;\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)\geq 0,$
@@
where max can be used instead of sup, because $\Delta(\mathcal{Z})$ is compact. Furthermore, $L_{\Omega}(\eta;q)$ is zero if and only if $q=\arg\max_{q\in\Delta(\mathcal{Z})}\mathbb{E}_{Z\sim q}[\eta(Z)]-\Omega(q)$. This maximization defines the $\Omega$-regularized prediction map $\Pi_{\Omega}:\mathcal{F}\rightarrow\Delta(\mathcal{Z})$,
@@
where $\mathcal{F}\subseteq\mathbb{R}^{\mathcal{Z}}$ is the set of functions for which the maximizer exists and is unique [10]. A trivial but important property of $\Pi_{\Omega}$ is invariance under global shifts of the argument: $\Pi_{\Omega}[\eta+C]=\Pi_{\Omega}[\eta]$, for any constant $C.$
@@
A particularly relevant choice for $\Omega$, which we will adopt in this paper, is the Tsallis negentropy [53],
$\Omega_{\rho}(q):=\left\{\begin{array}[]{ll}\displaystyle\mathbb{E}_{Z\sim q}[q(Z)^{\rho-1}-1]/\bigl{(}\rho(\rho-1)\bigr{)},&\text{if $\rho\neq 1$,}\\ \mathbb{E}_{Z\sim q}[\log q(Z)],&\text{if $\rho=1$,}\end{array}\right$. (10)
where $\rho>0$ is a parameter, called “entropic index”. Tsallis negentropies are continuous on $\rho$ and coincide with the Shannon negentropy when $\rho=1$. The $\Omega_{\rho}$-regularized prediction map $\Pi_{\Omega_{\rho}}$ is called $\rho$-entmax [47] and recovers softmax and sparsemax, for $\rho=1$ and $\rho=2$, respectively. For any $\rho>1$, $\Omega_{\rho}$-regularized prediction map will be sparse, with higher values of $\rho$ giving distributions with increasingly sparse supports.
Appendix A describes the relationship between FY losses based on Tsallis negentropies and the $\alpha$- and $q$-divergences of Amari and Ohara [4]. Amari’s $\alpha$-divergences, which have connections with Tsallis statistics, are related to but distinct from the divergences associated with Rényi entropies, which, somewhat confusingly, are also called $\alpha$-divergences; we do not address Variational Rényi approaches in this paper [36, 41, 34, 17, 18, 26].

### v3

@@ Abstract
Bayesian inference has a long history in machine learning but struggles in modern applications where exact inference is both infeasible and fragile due to the absence of well-specified priors and likelihoods. Variational inference with generalized objectives and approximating families aims to address these issues, but it is not always easy to adapt existing algorithms to the generalized setting. We present a new framework called Fenchel-Young variational inference (FYVI), which couples tightly the variational family to the objective. This coupling leads to easy and natural generalizations of existing variational methods to general posterior forms, which we exemplify through the expectation-maximization algorithm for latent variable models. We focus on a class of objectives based on Tsallis entropies which allow us to obtain generalized posteriors with support smaller than that of the prior. In several controlled experiments where modeling with sparse posteriors is intuitively appealing, we show that our generalized algorithms outperform established counterparts.
@@ 1 INTRODUCTION
Bayesian inference has long played a central role in machine learning, for example, as put forward in the seminal textbooks of bishop2006pattern and murphy2012machine. Despite its foundational role, the most modern applications in deep learning struggle to make use of Bayesian inference due to lack of well-specified priors, likelihoods, as well as the sheer scale of the problems making exact inference intractable. Pursuing early ideas of Zellner1988, recent work by knoblauch2022optimization and JMLR:v24:22-0291, as well as PAC-Bayesian approaches (see the recent survey of alquier2024user), advocate for an optimization-centric variational view of Bayesian inference. This view leads to models and algorithms with desirable Bayesian characteristics but which do not necessarily aim to execute or to approximate Bayes rule in a well-specified model.
JMLR:v24:22-0291 unify diverse, well-established Bayesian inference algorithms and modern deep learning methods under a common paradigm: Using natural-gradient descent to find exponential family posteriors over the unknown parameters or latent variables that strike a balance between a low expected risk and a small Kullback-Leibler divergence (KLD) from a given prior. This line of work led to recent successes, for instance, in Bayesian deep learning [vogn, ivon], large language model finetuning [cong2025improving], language generation [daheim2025uncertainty] and federated learning [mollenhoff2025federated]. However, it remains an open question how such formulations can be naturally extended from Kullback-Leibler divergences and exponential family posteriors to more general divergences and families which may work better in modern settings.
In this paper, we propose a new framework called Fenchel-Young variational inference (FYVI) which addresses this problem, and provides new generalized notions of FY free energy, FY evidence, FY evidence lower bound, and FY posterior. FYVI generalizes the formulations of [knoblauch2022optimization, JMLR:v24:22-0291], leading to a strictly larger class of statistical families which include generalized posteriors with limited (sparse) support, as well as heavy-tailed posteriors. A key property of the new framework is that it couples tightly the variational family to the variational objective, which we argue is the right modeling that leads to natural generalizations of existing algorithms to general posterior forms. This is where FYVI departs from knoblauch2022optimization who analyze arbitrary combinations of divergences, losses, and exponential variational families, and generalizes the modeling of JMLR:v24:22-0291 who are restricted to KLD and exponential families.
@@
The starting point for our new FYVI framework is the fact that the Bayesian posterior is the solution of an optimization over probability distributions where the objective is the expected log-likelihood plus the Kullback-Leibler divergence of a variational posterior from the prior [donsker1975asymptotic, csiszar1975divergence, Zellner1988]. The KLD formulation is equivalent to a maximum entropy problem that favors posteriors with high Shannon entropy [Jordan1999], giving a privileged position to exponential families [Grunwald_Dawid], following from the well-known fact that these are maximum Shannon entropy distributions under linear expectation constraints. The main idea of FYVI is to consider alternative optimality criteria that exploit the class of Fenchel-Young (FY) expressions [Blondel2020, martins2022sparse] and thereby extend the variational formulations of generalized posterior inference [bissiri2016general, knoblauch2022optimization, JMLR:v24:22-0291]. We remark that while Fenchel-Young losses have been used in the past for empirical risk minimization and discriminative learning [Blondel2020, martins2022sparse], our FYVI framework is radically different, as it uses FY for generalized Bayesian learning.
If we replace the KLD with an FY term in variational objectives, we obtain generalized deformed exponential family posteriors, which unlike exponential families may have variable support. We will focus on the particular case of Tsallis entropy FY objectives, which subsume the KLD. In particular, we consider sparse posteriors, but note that heavy-tailed posteriors are an important alternative case that also falls under the proposed framework. We derive a new EM-style [dempster1977maximum] algorithm called FYEM for parameter estimation, and describe gradient backpropagation for amortized inference with deformed exponential family posteriors. Either of these estimation methods could be used to fit arbitrary generative models specifiable as instances of FYVI, and the proposed algorithms are natural generalizations of well-known KLD-based counterparts.
@@
We experiment with the FYEM algorithm applied to Gaussian mixture model (GMM) [murphy2022probabilistic] estimation, and two FY variational autoencoders (VAEs) [Kingma2014, rezende2014stochastic], one applied to images and the other to documents. The FYEM algorithm for GMMs recovers both classical EM and hard EM as special cases and yields new generalizations, for instance, an adaptively sparse E-step. In the image FYVAE we obtain the best performance when we use a sparse posterior, as also observed in the concurrent work [qin2024kernel]. We also study the effect of changing the posterior form on a neural variational document model (NVDM) [miao2016neural]. All our experiments suggest that our generalized algorithms can outperform established counterparts.
@@ 2.1 Variational View of Bayesian Inference
Standard Bayesian Inference.
The Bayesian posterior $p_{Z|X}$ can be expressed in variational form as the distribution in $\Delta(\mathcal{Z})$ which strikes a balance between a small KLD from the prior $p_{Z}$ and a high expected log-likelihood $\log p_{X|Z}$, given some observation $x\in\mathcal{X}$ [Thanh2021, csiszar1975divergence, donsker1975asymptotic, Zellner1988]. Formally,
${p_{Z|X}(\cdot|x)}$
$\quad=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\,\underset{z\sim q}{\mathbb{E}}\bigl[\log p_{X|Z}(x|z)\bigr]\!-\!D_{\textnormal{KL}}(q\|p_{Z})$ (1)
$\quad=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\,\underset{z\sim q}{\mathbb{E}}\bigl[\log p_{X,Z}(x,z)\bigr]\!+\!H(q).$ (2)
The optimum value of (2), attained by $q=p_{Z|X}(\cdot|x)$, is equal to the log-evidence:
$\mathbb{E}_{z\sim p_{Z|X}(\cdot|x)}\bigl[\log p_{X,Z}(x,z)\bigr]+H\bigl(p_{Z|X}(\cdot|x)\bigr)$ (3)
$\qquad=\log p_{X}(x).$
More often than not the posterior and evidence cannot be easily obtained, which has stimulated decades of work on approximate Bayesian methods.
Variational Inference.
The goal of variational inference (VI) [Blei2017, Jordan1999] is to approximate the posterior and bound the evidence when these are intractable, by confining $q$ to a tractable family: $\mathcal{Q}\subseteq\Delta(\mathcal{Z})$. The familiar VI problem is
@@
and its optimal value is the evidence lower bound (ELBO): ${\mathbb{E}_{z\sim q_{x}^{\star}}}\bigl[\log p_{X,Z}(x,z)\bigr]+H\bigl(q_{x}^{\star}\bigr)\leq\log p_{X}(x)$. The negative of (4) is called the variational free energy.
Generalized VI.
@@
JMLR:v24:22-0291 extend (4) to loss-based posteriors using bissiri2016general’s generalization of classical VI in order to derive as special cases much of the core catalog of modern machine learning algorithms and to propose new ones. The extension consists of replacing the negative log-likelihood with a loss function $\ell(x;z)$:
$q_{x}^{\star}$ $=\arg\min_{q\in\mathcal{Q}}\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)\right]+D_{\text{KL}}(q\|p_{Z})$ (5)
$=\arg\min_{q\in\mathcal{Q}}\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)-\log p_{Z}(z)\right]-H(q),$ (6)
@@
where the notation $q_{x}^{\star}$ highlights the dependence on the observations $x$. JMLR:v24:22-0291 restrict $\mathcal{Q}$ to be a regular minimal exponential family, that is, $\mathcal{Q}=\{q(z)=h(z)\exp\bigl(\langle\lambda,t(z)\rangle-A(\lambda)\bigr),\mbox{for }\lambda\in\Lambda\}$, where $\Lambda$ is the parameter space, $t$ is a canonical sufficient statistic, $h$ is the base measure (independent of $\lambda$), and $A(\lambda)$ is the log-partition function. In a similar vein, knoblauch2022optimization propose generalized VI (GVI), by additionally considering divergences other than KLD, and introducing the rule of three, which refers to the choice of the three elements required for GVI: A loss function, a divergence measuring the deviation of variational posterior from prior, and a set of permissible solutions $\mathcal{Q}$.
2.2 Fenchel-Young Losses
@@
In a previously seemingly unrelated line of research, Blondel2020 propose Fenchel-Young (FY) losses, which generalize the ubiquitous cross-entropy loss, and can be seen as mixed space Bregman divergences (see also [amari2010information, amari2011geometry]). FY losses have been used for supervised problems, but not in any kind of Bayesian contexts like VI or latent variable models. Here we use FY expressions in a fundamentally different way since we do not use them to evaluate discrepancy between the score of some distribution and data. Our notation below treats both discrete and continuous domains in a unified way.
@@
Let $\Omega:\Delta(\mathcal{Z})\rightarrow\bar{\mathbb{R}}$ be a lower semi-continuous (l.s.c.) proper convex function, and let $\Omega^{*}$ be its Legendre-Fenchel conjugate [^1] Given a function $f:\mathcal{U}\rightarrow\bar{\mathbb{R}}$, its Fenchel conjugate $f^{*}:\mathcal{U}^{*}\rightarrow\bar{\mathbb{R}}$ is defined as $f^{*}(x)=\sup_{u\in\mathcal{U}}(\langle x,u\rangle-f(u)),$ where $\mathcal{U}^{*}$ is the dual space of $\mathcal{U}$. If $f$ is convex l.s.c., biconjugation holds: $f^{**}=f$. Assuming differentiability and that the supremum is attained, the maximizing argument in the definition of $f^{*}$ is $\arg\max_{u\in\mathcal{U}}(\langle x,u\rangle-f(u))=\nabla f^{*}(x)$. [Borwein, Bauschke]. And let $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ be a score function that assigns a real-valued score to each element of $\mathcal{Z}$. From a Bayesian perspective $\eta$ can be viewed as a log-prior. For $q\in\Delta(\mathcal{Z})$, over discrete or continuous $\mathcal{Z}$, the FY loss induced by $\Omega$ is a measure of how incompatible $\eta$ and $q$ are:
$D_{\Omega}(\eta;q)$ $=\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)$ (7)
$=\Omega^{*}(\eta)-\mathbb{E}_{z\sim q}[\eta(z)]+\Omega(q),$ (8)
@@
where the second equality is using the definition of the dual pairing $\langle q,\eta\rangle=\mathbb{E}_{z\sim q}[\eta(z)]$. When $\Omega$ is the Shannon negentropy $D_{\Omega}(\eta;q)=D_{\mathrm{KL}}(q||\exp(\eta)/\int\exp(\eta(z))dq(z))$. An FY loss is always non-negative, due to the Fenchel-Young inequality [Bauschke],
$\Omega^{*}(\eta)$ $=\sup_{q\in\Delta(\mathcal{Z})}\,\langle q,\eta\rangle-\Omega(q)\;$ (9)
$\Rightarrow\;\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)\geq 0.$ (10)
@@
Furthermore, $D_{\Omega}(\eta;q)$ is zero if and only if $q=\arg\max_{q\in\Delta(\mathcal{Z})}\mathbb{E}_{z\sim q}[\eta(z)]-\Omega(q)$. This defines the $\Omega$-regularized prediction map $\Pi_{\Omega}:\mathcal{F}\rightarrow\Delta(\mathcal{Z})$,
@@
where $\mathcal{F}\subseteq\mathbb{R}^{\mathcal{Z}}$ is the set of functions for which the maximizer exists and is unique [Blondel2020], and we exclude $\Omega$ for which $\mathcal{F}$ is the empty set. A trivial but important property of $\Pi_{\Omega}$ is invariance under global shifts of the argument: $\Pi_{\Omega}[\eta+C]=\Pi_{\Omega}[\eta]$, for any constant $C.$
@@
We choose $\Omega$ to be the Tsallis negentropy [tsallis2009introduction],
$\Omega_{\rho}(q)=\left\{\begin{array}[]{ll}\hskip-4.26773pt\mathbb{E}_{z\sim q}[q(z)^{\rho-1}-1]/\bigl(\rho^{2}-\rho\bigr),&\text{$\rho\neq 1$,}\\ \hskip-4.26773pt\mathbb{E}_{z\sim q}[\log q(z)],&\text{$\rho=1$,}\end{array}\right$. (14)
where $\rho>0$ is a parameter called the entropic index. Tsallis negentropies are continuous in $\rho$ and coincide with the Shannon negentropy when $\rho=1$. For finite $\mathcal{Z}$ the $\Omega_{\rho}$-regularized prediction map $\Pi_{\Omega_{\rho}}$ is called $\rho$-entmax [peters2019sparse] and recovers softmax and sparsemax, for $\rho=1$ and $\rho=2$ respectively. For any $\rho>1$, the $\Omega_{\rho}$-regularized prediction map will be sparse, with higher values of $\rho$ giving distributions with increasingly sparse supports.
In Appendix A we describe the relationship between Tsallis negentropy FY losses and the $\alpha$- and $q$-divergences of amari2011geometry. These $\alpha$-divergences are connected to Tsallis statistics, and are related to but distinct from the divergences associated with Rényi entropies, which, somewhat confusingly, are also often called $\alpha$-divergences; we do not address Variational Rényi approaches in this paper [li2016renyi, minka2005divergence, knoblauch2022optimization, daudel2023alpha, daudel2023monotonic, guilmeau2023variational].